OpenStudy (anonymous):

What is a formula to calculate area for a circle with diameter forty-six miles?

OpenStudy (jdoe0001):

area of a circle = \(\bf \large \pi r^2\)

OpenStudy (anonymous):

\[A = \pi r^{2}\] and \[r = \frac{ diameter }{ 2 }\]

OpenStudy (anonymous):

For this diameter I do not have the formula but, in general, for any diameter D, the solution is: PI x D^2/4

OpenStudy (anonymous):

ty all o_o